OpenStudy (anonymous):

Among all positive numbers a, b whose sum is 6 find those for which the product of the two numbers and their difference is largest.

OpenStudy (anonymous):

a + b = 6 a = 6 - b let y = maximum ab = y (6-b)b = y -b^2 + 6b = y [d/dx] -b^2 + 6b = [d/dx]y -2b + 6 = dy/dx dy/dx = 0 -2b +6 = 0 -2(b -3) = 0 Therefore, b = 3 a + b = 6 a + 3 = 6 a = 3 Therefore, b = 3

OpenStudy (anonymous):

Sorry but apparently this is not right because this is my online homework and its telling me its incorrect >_< help....